OpenStudy (anonymous):

Use two variables and two equations. Solve by any method you choose. Write your answer in the appropriate form. A lady has 11 animals. Some are cats and some are birds. If there are 34 legs among all her animals, how many animals does she have?

OpenStudy (anonymous):

@reach.koe

OpenStudy (austinl):

I am unsure how you would make two problems about this...

OpenStudy (anonymous):

kk thx anyway

OpenStudy (austinl):

Wait nvm.

OpenStudy (austinl):

x+y=11 4x+2y=34

OpenStudy (anonymous):

Wait a min. Do you want the number of animals (which is already given - 11) or number of cats & birds??? Assuming you want number of cats & birds. x = cats y = birds Solve for 4x + 2y = 34 (4 legs per cat & 2 legs per bird. 34 in all) x + y = 11 (total no of pets = 11) Solve the simultaneous equations ;)